Two trains leave the station at the same time, one heading west and the other east. THe westbound train travels 14 miles per hour faster than the eastbound train. If the two trains are 760 miles apart after 4 hours, what is the rate of the westbound train?
DO NOT ROUND

Answer by rfer(16322)

760/4=190
x+x+14=190
2x=176
x=88 mph
x+14=102 mph